# HueManity: Probing Fine-Grained Visual Perception in MLLMs

### v2

@@ Abstract
Multimodal Large Language Models (MLLMs) excel at high-level visual reasoning, but their performance on nuanced perceptual tasks remains surprisingly limited. We present HueManity, a benchmark designed to assess visual perception in MLLMs. The dataset comprises 83,850 images featuring two-character alphanumeric strings embedded in Ishihara test style dot patterns, challenging models on precise pattern recognition. Our evaluation of nine state-of-the-art MLLMs on HueManity demonstrates a significant performance deficit compared to human and traditional computer vision baselines. The best-performing MLLM achieved a 33.6% accuracy on the numeric ‘easy’ task and a striking 3% on the alphanumeric ‘hard’ task. In contrast, human participants achieved near-perfect scores (100% and 95.6%) [^1] Human evaluations utilized 100-image representative subsets, sampled from the model evaluation sets for each respective task., and a fine-tuned ResNet50 model reached accuracies of 96.5% and 94.5%. These results highlight a critical gap in the visual capabilities of current MLLMs. Our analysis further explores potential architectural and training-paradigm factors contributing to this perceptual gap in MLLMs. We open-source HueManity dataset and code to foster further research in improving perceptual robustness of MLLMs.
Code: https://github.com/rynaa/huemanity Dataset: https://huggingface.co/datasets/Jayant-Sravan/HueManity
HueManity: Probing Fine-Grained Visual Perception in MLLMs
Rynaa Groverthanks: Equal Contribution Google rynaa@google.com                                   Jayant Sravan Tamarapalli11footnotemark: 1 Google jayantsravan@google.com
Sahiti Yerramilli11footnotemark: 1 Google sahitiy@google.com                                   Nilay Pande11footnotemark: 1 Waymo nilayp@waymo.com
Figure 1: HueManity — A new benchmark for MLLM fine-grained visual perception. The pipeline creates a character mask from alphanumeric characters and renders it as an Ishihara-style pattern. While models achieve high accuracy on clear masks, they struggle significantly on the challenging pattern images.
@@ 1 Introduction
The trajectory of Multimodal Large Language Models (MLLMs) Team et al. (2023); Achiam et al. (2023); Bai et al. (2023a); Li et al. (2023b); Gong et al. (2023); Liu et al. (2024a, 2023); Anthropic (2025) in recent years has been marked by impressive advancements, demonstrating sophisticated capabilities in bridging visual and textual information. Their capabilities extend well beyond simple image labeling Russakovsky et al. (2015); Deng (2012), enabling complex tasks like generating detailed image descriptions Dong et al. (2024); Fu et al. (2024a), answering intricate visual questions requiring inference about relationships and activities Weng et al. (2025); Chen et al. (2025); Kuang et al. (2024), and participating in nuanced dialogue about visual content Cao et al. (2024). A key factor in their success is pre-training on vast web-scale image-text datasets, which facilitates learning powerful representations that capture high-level semantic links between visual features and language Jia et al. (2021); Radford et al. (2021); Schuhmann et al. (2022); Alayrac et al. (2022); Qi et al. (2020); Zhai et al. (2022); Pham et al. (2023). Consequently, MLLMs perform strongly on tasks benefiting from this conceptual understanding, including recognizing common objects, interpreting general scene structure, and relating them to text.
@@
However, the predominant evaluation paradigms for MLLMs have centered on these conceptual capabilities, largely overlooking their fine-grained perceptual acuity Bai et al. (2023b); Li et al. (2024, 2023a); Xu et al. (2024); Yin et al. (2023); Liu et al. (2023). This leaves a critical gap in understanding their performance on tasks demanding precise visual discernment – such as intricate pattern recognition, subtle feature differentiation, and robust segregation within visually cluttered background. Unlike tasks solvable through broad semantic association, these perceptual challenges require a more fundamental visual processing capacity, akin to the human ability to resolve intricate details. This paper introduces a benchmark specifically designed to probe this nuanced dimension of MLLM performance. Our methodology draws inspiration from the principles of Ishihara plates Clark (1924), a technique traditionally employed in human ophthalmology to assess color vision by embedding figures (like numbers or paths) within fields of multicolored, varied-size dots. It is crucial to clarify that HueManity does not aim to diagnose ‘color blindness’ in MLLMs. Instead, our Ishihara-style stimuli, created using controlled generation techniques, rigorously test an MLLM’s fundamental ability to identify embedded alphanumeric characters by their subtle color and luminance contrasts within visually cluttered dot patterns.
Success on the HueManity benchmark serves as a crucial indicator of an MLLM’s potential for robust visual understanding in complex, real-world scenarios. Unlike often curated benchmark datasets, real-world visual environments are frequently characterized by clutter, partial occlusions, variable lighting, and unconventional information presentation. The ability of an MLLM to reliably parse characters in our Ishihara-style plates is intended to assess its resilience to visual clutter and its pattern recognition capabilities — foundational skills often linked to dependable performance in challenging visual settings. Furthermore, mastering this level of fine-grained perception is foundational for more intricate forms of visual reasoning. An MLLM’s struggles with basic pattern recognition under challenging conditions could imply difficulties in performing higher-order tasks that depend on the accurate interpretation of subtle details. Thus, HueManity serves not merely as a test of pattern recognition, but as a probe into architectural or training limitations that hinder MLLMs from achieving comprehensive, detailed visual intelligence.
@@
To address this identified gap and facilitate further research in this domain, this paper makes the following specific contributions:
@@
    We conducted a comprehensive evaluation of nine state-of-the-art MLLMs on HueManity’s easy and hard sets. Findings reveal a significant performance gap compared to strong human and fine-tuned ResNet50 baselines. Baseline success suggests MLLM limitations are due to their current architectures or training for fine-grained perception, rather than the task being intractable.
@@
    We release open-source code for generating challenging Ishihara-style perceptual stimuli (alphanumeric characters embedded in dot patterns), enabling reproducible research and community-driven extensions.
@@ 2.1 Multimodal Models
With the remarkable advancements of Large Language Models (LLMs), recent research has extended their capabilities to multimodal domains by integrating visual information, giving rise to Multimodal Large Language Models (MLLMs) Team et al. (2023); Achiam et al. (2023); Bai et al. (2023a); Li et al. (2023b); Gong et al. (2023); Liu et al. (2024a, 2023). These models typically align visual features from pre-trained image encoders with LLMs via modality adaptation layers. Early works like BLIP-2 Li et al. (2023b) pioneered this architecture by first pre-training on image-text datasets and fine-tuning on task-specific benchmarks such as Visual Question Answering (VQA). Subsequent models like LLaVA Liu et al. (2023) advanced this approach by leveraging synthetic instruction-following data in VQA formats, significantly improving instruction tuning performance. More recent efforts have expanded into video understanding and even image generation, showcasing the versatility of MLLMs across modalities. However, much of their celebrated success in visual tasks often appears intertwined with, and perhaps reliant upon, their powerful language capabilities to interpret, reason about, and generate text from visual content. This inherent strength in linguistic processing may have led to less emphasis on developing or scrutinizing their more fundamental, non-linguistic visual perception skills. HueManity directly addresses this gap in scrutinizing MLLM visual perception. It is a benchmark specifically designed to isolate and rigorously evaluate these core abilities.
@@ 2.2 MLLM Evaluation
Multimodal Large Language Models (MLLMs) are highly capable of global image understanding and reasoning but consistently struggle with fine-grained visual tasks such as precise recognition and localization Huang and Zhang (2024); Li et al. (2024). Several benchmarks have been proposed to address these challenges. TouchStone Bai et al. (2023b) offers 908 manually annotated visual dialog questions across five abilities and 27 sub-tasks, while LLaVA-Bench Liu et al. (2023) includes 24 images with 60 curated questions covering diverse content like scenes, memes, and sketches. To automate evaluation, LLaVA-Bench Liu et al. (2023), LAMM Yin et al. (2023), and TouchStone Bai et al. (2023b) rely on GPT-based models to judge relevance and accuracy, but this introduces inherent reliability issues and cost inefficiencies. Similarly, LVLM-eHub Xu et al. (2024) aggregates multiple vision benchmarks but still depends on human annotators to compare model outputs, making it both subjective and expensive. More structured efforts like MME and MMBench Liu et al. (2024b) attempt to provide objective evaluations with multiple-choice questions across a wide range of ability dimensions. MME pioneered multimodal Yes/No questions for perception and reasoning, with MM-Vet Yu et al. (2023) and MMBench Liu et al. (2024b) extending coverage to sub-tasks like OCR, math, and recognition, yet they all heavily rely on existing VQA datasets or GPT-generated questions. SEED-Bench Li et al. (2024, 2023a) scales up to 24,000 human-annotated multiple-choice questions covering varied input-output modalities, but its challenges remain simple, with most open-source models reaching 30–60% accuracy at the easiest level. Blink Fu et al. (2024b) attempts a more holistic evaluation across 14 perception tasks with 3,900 questions and 7,300 images, but does not assess models on combinations of tasks, and remains less challenging, as evidenced by higher model accuracies.
HueManity uniquely evaluates Multimodal Large Language Models (MLLMs) on the foundational visual skill of discerning patterns from subtle cues in cluttered environments, employing alphanumeric characters within Ishihara-style dot patterns with controlled color contrasts as a robust proxy. This stimulus design, combined with procedural generation and exact-match evaluation, establishes a scalable, objective, and reliable methodology distinct from more subjective or resource-intensive techniques.
@@ 3 Data Creation
The HueManity dataset comprises 83,850 images, each pairing a two-character alphanumeric string with its ground truth label and full generation parameters. These strings are formed from lowercase letters (a-z), uppercase letters (A-Z), and digits (0-9). To enhance clarity and prevent evaluation errors, we excluded visually ambiguous characters (‘l’, ‘I’, ‘J’, ‘O’) and combinations commencing with ‘0’ (e.g., ‘01’, to avoid ‘1’ vs. ‘01’ prediction conflicts). The entire dataset was produced by rendering all valid two-character combinations using 25 meticulously curated color pairs, which then formed the basis for our evaluation sets.
The full HueManity dataset, encompassing all 83,850 generated images, serves as a comprehensive resource for evaluating MLLM perception. However, for the specific MLLM evaluations reported in this paper, particularly considering practical constraints such as time and API access costs, we defined two distinct tasks using carefully sampled subsets. Each subset comprises 1,000 images, randomly selected to be representative for these focused evaluations:
  1. 1.
    Number Recognition Set (Used for the ‘Easier’ Task): This subset of 1,000 images is used to evaluate models on the Number Identification task. It exclusively contains images displaying two-character strings composed solely of numeric digits (e.g., ‘17’, ‘83’, ‘05’).
  2. 2.
    Text Recognition Set (Used for the ‘Harder’ Task): This subset of 1,000 images, drawn from the broader pool of alphanumeric combinations, is employed for the Text Identification task. It features diverse two-character alphanumeric strings, including various combinations of letters and digits (e.g., ‘A7’, ‘b9’, ‘XG’).
3.1 Text Mask Generator
@@
The initial stage in our data pipeline is the creation of a 900x900 pixel resolution binary text mask for each two-character string. Using Pygame, we render these strings in white against a black background. To ensure both clear readability and adequate character width suitable for the subsequent dot-based rendering, all masks utilize the DejaVu Sans font (size 550, bold and italic styles), as exemplified in Figure 2(a).
(a) Binary Text Mask
(b) Ishihara pattern
Figure 2: Image generation process for the alphanumeric string “Y8”. (a) The binary text mask. (b) The corresponding Ishihara-style dot pattern.
3.2 Ishihara-Style Pattern Generation
Our pattern generator, adapted from an open-source Pygame project [^2] https://github.com/hakrackete/Ishihara-color-plate-generator, iteratively populates the image with non-overlapping circles. Over 30,000 iterations, the generator performs the following steps for each potential circle:
  1. 1.
    Placement: Randomly sample (x,y) coordinates and compute the maximum non-colliding radius (ranging from 4 to 15 pixels).
  2. 2.
    Initial Color: Assign initial color (foreground/background) based on whether the center falls within the character mask (Fig. 2(a)).
  3. 3.
    Color Transformation: Apply three randomized transformations to the initial color: gradient shift towards the other pair color, RGB color shift (range $[-30,+30]$), and RGB lightness scaling (factor 0.66-1.5).
  4. 4.
    Rendering: Render the transformed circle at the computed position.
This iterative process ultimately yields the final, dense Ishihara-style pattern (as shown in Figure 2(b)).
3.3 Color Pairs Selection
@@
The 25 distinct foreground-background color pairs used in the HueManity stimuli were meticulously selected through a multi-stage process to ensure a balance between perceptual challenge and unambiguous human legibility. This procedure involved quantitative CIEDE2000 Luo et al. (2001) analysis and extensive manual verification, with the full methodology detailed in Appendix B.
4 Experiments
@@
We evaluated a diverse set of nine state-of-the-art Multimodal Large Language Models (MLLMs), encompassing prominent commercial APIs and publicly available open-source models. Model inference was orchestrated using Promptfoo [^3] https://www.promptfoo.dev, a platform facilitating reproducible benchmarking through flexible prompt definition and API integration. Open-source models were hosted locally via Ollama [^4] https://ollama.com and inferred on a single NVIDIA A100 GPU.
Evaluation focused on two primary tasks: numerical recognition (digits only) and alphanumeric recognition (digits and letters). For each of these tasks, a subset of 1,000 points from HueManity: two-character strings and their corresponding generated images were used. Models were evaluated on two types of visual stimuli for each datapoint:
  • HueManity Plates: The 1,000 randomly sampled Ishihara-style dot pattern images from the HueManity benchmark.
  • Text Masks: The corresponding 1,000 binary text mask images (i.e., white text on a black background, as exemplified in Figure 2(a)). This additional evaluation on text masks was performed to establish a baseline for each model’s ability to recognize the characters in a clear, unobstructed format, helping to differentiate fundamental OCR capabilities from performance on the perceptually challenging dot patterns.
All images (both HueManity plates and text masks) were provided to the models encoded in Base64 format. The following task-specific prompts were used for both stimulus types:
  • Number Recognition Prompt: "What is the number in this image? Strictly stick to the format: Answer: [number in the image]"
@@
  • Text Recognition Prompt: "What is the exact text in this image? It has only alpha-numeric characters excluding small L, capital O, capital I, and capital J to avoid ambiguity. Strictly stick to the format: Answer: [exact text in the image]"
4.1 Human Performance Evaluations
To establish a human performance baseline, evaluations involved three adult volunteers, all with self-reported normal color vision. This group size was considered adequate because identifying characters in HueManity plates is a rudimentary perceptual task for humans with normal vision, one that requires no specialized skills and typically exhibits low inter-volunteer performance variance. For each of the two tasks (numerical and alphanumeric recognition), these volunteers were tested on a representative subset of 100 images, randomly sampled from the corresponding 1,000-image evaluation sets used for MLLMs.
Volunteers viewed the images at their original 900x900 pixel resolution within a Google Sheets document. Using the identical prompts provided to the MLLMs, they were asked to identify the embedded characters, entering their responses directly into the document. This methodology ensured a direct comparison of human and MLLM performance under equivalent conditions.
4.2 Traditional Computer Vision Baseline (ResNet50)
As a representative traditional computer vision baseline, we trained and evaluated a ResNet50 model. We utilized a ResNet50 pre-trained on ImageNet, obtained from the PyTorch vision library.
@@ 4.2 Traditional Computer Vision Baseline (ResNet50)
The standard classification layer of the ResNet50 was replaced with two independent classification heads. Each head was designed to predict one character, treating the task as two independent character recognition problems. For the purpose of fine-tuning this model, we utilized 2,000 images randomly sampled from the broader HueManity dataset, ensuring these were distinct from the final evaluation subsets. Training was conducted for 30 epochs using the Adam optimizer with a learning rate of $31e-$ 3. The loss function for training was the sum of the cross-entropy losses calculated independently for each of the two classification heads. The trained model was evaluated on the same 1,000-image subsets used for the MLLM evaluations.

### v5

@@ Abstract
Recent Multimodal Large Language Models (MLLMs) demonstrate strong high-level visual reasoning on tasks such as visual question answering and image captioning. Yet existing benchmarks largely overlook their ability to capture fine-grained perceptual details. As MLLMs are increasingly deployed in safety and reliability critical settings, perceptual acuity becomes essential. We present HueManity, a scalable automated benchmark for assessing fine-grained visual perception in MLLMs. HueManity comprises 83,850 Ishihara-style images embedding alphanumeric strings, designed to evaluate pattern recognition, a core aspect of visual understanding. Our evaluation of nine state-of-the-art MLLMs uncovers a striking performance deficit: the strongest model achieved only 33.6% accuracy on a simple numeric task and 3% on a harder alphanumeric task, compared to near-ceiling performance from humans (99.38%, 93.25%) and a fine-tuned ResNet-50 (96.5%, 94.5%). These findings expose a critical weakness in MLLMs’ perceptual grounding, one that remains obscured by conventional benchmarks emphasizing high-level semantics.
Code: https://github.com/rynaa/huemanity
Dataset: https://huggingface.co/datasets/Jayant-Sravan/HueManity
Figure 1: We present HueManity - an automated scalable benchmark for evaluation of fine-grained visual perception in MLLMs. The pipeline (left) embeds characters within challenging Ishihara-style image patterns, while ensuring human readability of the generated images. Experiments reveal (right) that humans and a fine-tuned ResNet50 baseline significantly outperform top-5 leading closed-source and open-source MLLMs, exposing a critical lack of fine-grained understanding.
@@ 1 Introduction
The trajectory of Multimodal Large Language Models (MLLMs) [13, 1, 5, 27, 19, 29, 4] has been marked by impressive advancements, demonstrating sophisticated capabilities in bridging visual and textual information. Their skills extend beyond simple image labeling [36, 14], enabling complex tasks like generating detailed image descriptions [15, 18], answering intricate visual questions requiring inference about relationships and activities [42, 10, 24], and participating in nuanced dialogue about visual content [9]. This success is largely attributed to pre-training on vast, web-scale image-text datasets, which has cultivated a powerful ability to map high-level semantic concepts between modalities [23, 34, 38, 3, 33, 46, 32].
However, this focus on semantic understanding, both in training and in predominant evaluation paradigms [7, 25, 26, 43, 45, 29], has created a critical blind spot: the models’ foundational perceptual acuity remains largely unprobed. Human vision is not merely a semantic engine, it is fundamentally a system for perceptual organization, excelling at extracting coherent signals from visually cluttered environments. This ability to discern patterns from subtle, low-level cues — like color, luminance, and texture — is a prerequisite for virtually all higher-level visual reasoning.
@@
Existing benchmarks have predominantly centered on the conceptual capabilities of MLLMs, leaving their resilience to perceptual challenges like pattern recognition and feature differentiation in cluttered scenes largely unevaluated. This paper introduces HueManity, a benchmark specifically designed to probe this gap. Our methodology is inspired by Ishihara plates [12], a classic tool from human ophthalmology designed to isolate a specific perceptual skill: figure-ground segregation based on subtle color cues. It is crucial to clarify that HueManity does not aim to diagnose ‘color blindness’ in MLLMs. Instead, our Ishihara-style stimuli, created using controlled generation techniques, test an MLLM’s fundamental ability to identify embedded alphanumeric characters by their subtle color and luminance contrasts within visually cluttered dot patterns.
The ability to parse characters from our Ishihara-style plates is a direct proxy for an MLLM’s capacity to handle real-world visual challenges characterized by clutter, partial occlusions, and variable lighting. An MLLM that cannot distinguish a number embedded in a field of dots may likewise struggle to perform optical character recognition (OCR) on text seen through a rainy window or accurately transcribe text from a crumpled, faded receipt. Therefore, HueManity serves not merely as a test of pattern recognition, but as a diagnostic tool to probe the limitations that prevent MLLMs from achieving robust, human-like visual intelligence. Our findings reveal that this novel evaluation exposes a profound deficit in current models, a result with significant implications for deploying MLLMs in domains requiring human-like perceptual understanding, from autonomous systems to medical imaging [44].
@@
To address this identified gap and facilitate further research in this domain, this paper makes the following contributions:
@@
    We conduct a comprehensive evaluation of nine state-of-the-art MLLMs, revealing a significant performance gap when compared to strong human and fine-tuned ResNet50 baselines. This suggests MLLM limitations are architectural rather than the task being intractable
@@
    We release open-source code for generating challenging Ishihara-style perceptual stimuli, enabling reproducible research and community-driven extensions.
@@ 2.1 Multimodal Models
With the remarkable advancements of Large Language Models (LLMs), recent research has extended their capabilities to multimodal domains by integrating visual information, giving rise to Multimodal Large Language Models (MLLMs) [13, 1, 5, 27, 19, 29]. These models typically align visual features from pre-trained image encoders with LLMs via modality adaptation layers. Early works like BLIP-2 [27] pioneered this architecture by first pre-training on image-text datasets and fine-tuning on task-specific benchmarks such as Visual Question Answering (VQA). Subsequent models like LLaVA [29] advanced this approach by leveraging synthetic instruction-following data in VQA formats, significantly improving instruction tuning performance. More recent efforts have expanded into video understanding and even image generation [8, 37], showcasing the versatility of MLLMs across modalities. However, this celebrated success in visual tasks often appears reliant on their powerful language capabilities for reasoning and interpretation, potentially overshadowing the need to scrutinize their fundamental perception skills. Addressing this gap, HueManity is a benchmark designed specifically to isolate and probe these visual abilities.
@@ 2.2 MLLM Evaluation
While MLLMs excel at global image understanding, they often struggle with fine-grained tasks requiring precise recognition and localization [22, 25]. Existing benchmarks designed to probe these limitations face three primary methodological challenges:
Semantic Entanglement:In many perception benchmarks, like Eyes Wide Shut [40], MERLIM [41]), perception is often “entangled” with reasoning. Questions focusing on semantic concepts like “left/right” or “yellow animal” allow models to rely on language priors rather than pure visual input.
Evaluation: Early benchmarks like LLaVA-Bench [29], and TouchStone [7] rely on GPT-based judges. However, LLM-based evaluation is unreliable for fine-grained perception since the best models are also not very adept. Other efforts, such as LVLM-eHub [43], depend on expensive human annotators, limiting scalability.
Data Integrity and Scale: To improve reliability, MME [17], MMBench [30], and SEED-Bench [25] adopted objective, multiple-choice formats. However, these often draw from existing VQA datasets, raising concerns about data contamination. While ZeroBench [35] offers increased difficulty, its utility for fine-grained perception is limited by its broad focus and small scale. Similarly, illusion-based benchmarks [20, 39] are often constrained by small dataset sizes and a reliance on common internet images.
HueManity departs from these approaches by explicitly disentangling perception from semantics. Unlike benchmarks where linguistic context could help guess an answer, HueManity provides no semantic cues to predict a hidden digit like ‘26’. The model must perceive the underlying dot patterns directly, making it impossible to rely on language priors. Through procedural generation and an exact-match evaluation framework, HueManity offers a novel, scalable, and entirely objective methodology for evaluating a core perceptual capability that remains a challenge for state-of-the-art MLLMs.
Table 1: Comparison of MLLM evaluation benchmarks across key methodological attributes. Novelty refers to the use of non-internet, generated images.
Benchmark Data Size Automatic Annotation Novel Images Answer Form Automated Evaluation
LVLM-eHub - Free-form
MMBench 3,217 Multi-choice
SEED-Bench2 24,371 Multi-choice
Blink 3,807 Multi-choice
ZeroBench 100 Free-form
HueManity 83,850 Exact Match
@@ 3 Data Creation
The HueManity benchmark is built upon a dataset of 83,850 images, each featuring a two-character alphanumeric string rendered as an Ishihara-style dot pattern, along with its corresponding ground truth label and generation parameters. Our generation process involves creating a base text mask, selecting a color palette, and rendering the final dot pattern. Figure 1 illustrates the core stages of this pipeline.
@@
Text Mask Generator. The first step is to create a 900x900 pixel binary mask for each two-character string. We use the Pygame library to render the string in white on a black background. To ensure the characters are thick enough for dot-based rendering and remain clearly legible, we use the DejaVu Sans font, styled in bold and italic at a size of 550 (Figure 1).
@@
Color Pairs Selection To generate the patterns, we meticulously selected 25 distinct foreground-background color pairs. This selection process involved a multi-stage procedure combining quantitative analysis using the CIEDE2000 color-difference formula [31] with extensive manual verification. This ensures that the color pairs are balanced for perceptual difficulty while remaining legible to humans (see Appendix D for details).
Ishihara-Style Pattern Generation. Our pattern generator, adapted from an open-source Pygame project [^1] https://github.com/hakrackete/Ishihara-color-plate-generator, iteratively populates the image with non-overlapping circles. Over 30,000 iterations, the generator randomly places circles, computes their maximum non-colliding radius (4-15 pixels), and assigns it a color. The color assignment depends on whether the circle’s center falls within the character region of the text mask. This initial color then undergoes three randomized transformations: a gradient shift towards the other color, an RGB color shift (range [-30, +30]), and an RGB lightness scaling (factor 0.66 - 1.5). These transformed circles are then rendered, resulting in the final dense Ishihara-style pattern (Figure 2).
Figure 2: Qualitative examples showing predictions of 4 representative MLLMs vs baselines on numeric and alphanumeric tasks.
4 Experimental Setup
Benchmark Tasks and Evaluation Sets. From the generated stimuli, we designed a benchmark with two recognition tasks of varying difficulty and two visual conditions to isolate specific model capabilities. There are two distinct recognition tasks of varying difficulty: a numeric task with a label space of 90 possible outputs and a more challenging alphanumeric task with 3,364 possible outputs (Refer to Appendix B.1). The larger search space makes the alphanumeric task inherently more difficult.
  • $\circ$
    Number Recognition Set (Easier Task): This subset contains images with two-digit numeric strings (e.g., 17, 83, 65). We exclude leading-zero numbers (‘00’-‘09’) to prevent ambiguity between single and double-digit interpretations.
  • $\circ$
    Alphanumeric Recognition Set (Harder Task): This subset includes two-character strings (e.g., A7, 9b, XG) formed from lowercase letters (a-z), uppercase letters (A-Z), and digits (0-9). We excluded visually ambiguous characters (‘l’, ‘I’, ‘J’, ‘O’) from the character space.
Further, to disentangle the challenge of perceptual grouping from fundamental character recognition, we evaluate models under two visual conditions for each image. Due to computational and API cost constraints, our MLLM evaluations were conducted on subsets of 1,000 images, randomly sampled for both the numeric and alphanumeric tasks.
  • $\circ$
    Ishihara Pattern: Primary task using the Ishihara-style dot pattern images.
  • $\circ$
    Text Masks: Control task using the binary text mask used in the creation of Ishihara Pattern images. This condition establishes a baseline for each model’s fundamental OCR capability, helping to disentangle it from performance on the more perceptually complex patterns.
@@ 4 Experimental Setup
Multimodal Large Language Models (MLLMs). We evaluated a diverse set of nine Multimodal Large Language Models (MLLMs), including both commercial API-based and publicly available open-source models. Model inference was orchestrated using Promptfoo [^2] https://www.promptfoo.dev, a platform facilitating reproducible benchmarking through flexible prompt definition and API integration. Open-source models were hosted locally via Ollama [^3] https://ollama.com and inferred on a single NVIDIA A100 GPU. Images were Base64 encoded at 900x900 original resolution and submitted with the following task-specific prompts, which were kept consistent across all models and conditions:
@@
Number Recognition Prompt: “What is the number in this image? Strictly stick to the format: Answer: [number in the image]” Text Recognition Prompt: “What is the exact text in this image? It has only alpha-numeric characters excluding small l, capital O, capital I, and capital J to avoid ambiguity. Strictly stick to the format: Answer: [exact text in the image]”
Human Performance Baseline. To establish a human performance baseline, we recruited 16 adult volunteers from diverse age groups with self-reported normal color vision. Each participant was evaluated on 100 Ishihara pattern images for each of the tasks (numerical and alphanumeric recognition). These images were randomly sampled from the 1,000-image sets used in the MLLM evaluations. Participants viewed the 900x900 pixel images in a Google Sheets document and provided their responses to the same prompts given to the models, ensuring a direct and methodologically consistent comparison.
@@
Traditional Computer Vision Baseline (ResNet50). As a representative traditional computer vision baseline, we trained and evaluated a ResNet50 model. We utilized a ResNet50 pre-trained on ImageNet, obtained from the PyTorch vision library. The standard classification layer of the ResNet50 was replaced with two independent classification heads. Each head was designed to predict one character, treating the task as two independent character recognition problems. For the purpose of fine-tuning this model, we utilized 2,000 images randomly sampled from the broader HueManity dataset, ensuring these were distinct from the final evaluation subsets. Training was conducted for 30 epochs using the Adam optimizer with a learning rate of 1e31e-3. The loss function was the sum of the cross-entropy losses calculated independently for each of the two classification heads. The trained model was evaluated on the same 1,000-image subsets used for the MLLM evaluations.